# CleanMel: Mel-Spectrogram Enhancement for Improving Both Speech Quality and ASR

### v1

@@ III-B2 Cross-band block
The cross-band block is used to learn full-band/cross-band dependencies of signals. In the original SpatialNet [21], the cross-band block is designed for learning the linear relationship of inter-channel features (e.g. inter-channel phase different) across frequencies. In this work, single channel does not have such inter-channel information. Instead, the cross-band block can learn the full-band spectral pattern of signals in the linear/Mel-frequency domain, which is also critical information for (especially single-channel) speech enhancement. The cross-band block processes frames independently along the frequency dimension, and all frames share the same network.
Specifically, we adopt the original cross-band block as in SpatialNet, which is composed of cascaded frequency convolutional layer (F-GConv1d), across-frequency linear layers (F-Linear) and a second frequency convolution layer. The frequency convolutional layers perform 1-D convolution along frequency to learn correlations between adjacent frequencies. The across-frequency linear layer processes all frequencies together to learn full-band dependencies. One cross-band block is first applied in the linear-frequency domain to learn detailed full-band correlations. To reduce the model complexity, the hidden dimension $H$ is compressed to a much smaller dimension, such as $H/12$, and then each hidden dimension is independently processed by the across-frequency linear layer with a complexity of $F^{2}$. After Mel-filterbank, the cross-band blocks learn full-band correlations across Mel frequencies, where the model complexity is largely reduced from $F^{2}$ to $F_{\text{mel}}^{2}$. Correspondingly, the hidden dimension is remained as $H$ (no longer compressed) to reinforce the capability of full-band learning. We set all the Mel-frequency cross-band blocks to share the same across-frequency linear layers.
@@ III-C Back-ends
At the inference stage, CleanMel is followed by either an ASR model or a neural vocoder. The ASR model and neural vocoder are separately trained with the CleanMel network.
@@ III-C1 ASR
At inference, the enhanced logMel-spectrogram is directly fed to an already-trained ASR system, without performing any fine-tuning or joint-training. Different ASR systems may have different configurations in STFT settings, number of mel frequencies and base of logarithm. To seamlessly integrate the enhanced logMel-spectrogram into one ASR model, our CleanMel would adopt the same configurations as the ASR model. The training cost of CleanMel is not very high, so it can be easily re-trained for a new ASR system, especially for those large-scale ASR systems and already-deployed ASR systems. In this work, we only conduct offline ASR combined with offline CleanMel.
@@ III-C2 Neural vocoder
The vocoder we adopt in this work is Vocos [26], a recently proposed Generative Adversarial Network (GAN)-based neural vocoder. The generator of Vocos predicts the STFT coefficients of speech at frame level and then generates waveform through inverse STFT. Vocos uses the multiple-discriminators and multiple-losses proposed in HiFi-GAN [27], but it significantly improves the computational efficiency compared to HiFi-GAN that directly generates waveform at sample level. In this work, to unify the front-end and back-end processing, we have made several necessary modifications to Vocos as follows:
  • The magnitude-based Mel-spectrogram of original Vocos is modified as power-based to be consistent with the front-end and ASR models, where the two cases were shown to achieve similar performance in our preliminary experiments.
  • The sampling rate of signals, the STFT configurations and the number of Mel-frequencies of the Vocos are adjusted according to the setup of front-end and ASR models.
  • The original Vocos is designed for offline processing, as it employs non-causal convolution layers. To enable online processing, we modified Vocos to be causal by substituting each non-causal convolution layer with its causal version. The online vocoder still performs quite well. In addition, to reduce the computational complexity of online processing, the 75% STFT overlap of original Vocos is reduced to be 50% overlap, which still achieves comparable performance.
The offline and online Vocos are used to work with the offline and online CleanMel, respectively. Vocos models are trained with our direct-path target speech $x(n)$.
@@ III-D Signal Normalization
When separate front-end and back-end models are cascaded, signal normalization should be performed not only to facilitate the training of respective models but also to align the signal level of cascaded models. For offline processing, the normalization method used in Vocos is also applied for CleanMel. Specifically, a random gain is applied to time domain signal to ensure that the maximum level of the resulting signal lies within the range of -1 to -6 dBFS. This normalization manner ensures the maximum level of sample values is close to and smaller than 1, thence the generated waveform can be directly played with full volume and without clipping effect. For CleanMel, we apply this normalization to noisy signal and utilize the same gain of noisy signal to the corresponding clean target signal. In this way, the enhanced Mel-spectrogram can be directly fed to Vocos. When applying this time-domain normalization, the clip value $\epsilon$ for computing logMel-spectrogram is set to 1e-5 for both CleanMel and Vocos. As for ASR, ASR models normally have a separate Mel-spectrogram normalization operation, which will be applied to re-normalize the enhanced Mel-spectrogram.
For online processing, the time domain normalization method is no longer applicable. Instead, an online STFT-domain normalization is used for both CleanMel and Vocos. Specifically, for CleanMel, the noisy and target speech are normalized in the STFT domain as $\tilde{Y}(f,t)=Y(f,t)/\mu(t)$ and $\tilde{X}(f,t)=X(f,t)/\mu(t)$, where $\mu(t)$ is a recursively calculated meanvalue of STFT magnitude of $Y(f,t)$, i.e. $\mu(t)=\alpha\mu(t-1)+(1-\alpha)\frac{1}{F}\sum_{f=0}^{F-1}|Y(f,t)|$. The smoothing weight is set to $\alpha=\frac{K-1}{K+1}$, by which the recursive smoothing is equivalent to using a $K$-long rectangle smoothing window. When training Vocos using target speech signal $x(n)$, we still first apply the time domain normalization mentioned above, then apply an extra online normalization. Specifically, $\mu(t)$ is computed with and also applied to $X(f,t)$, and then the corresponding logMel-spectrogram is computed as the input of Vocos generator. Accordingly, the output of Vocos generator (before applying inverse STFT) would be an estimation of normalized $X(f,t)$. To go back to the signal level of time domain normalization, the recursive normalization factor $\mu(t)$ is multiplied back to the estimation of normalized $X(f,t)$ and then applying inverse STFT, after which Vocos losses (including Mel loss and discriminator losses) are computed. At inference, online normalization is applied to the noisy input, and the enhanced logMel-spectrogram is directly fed into Vocos. The recursive normalization factor computed with the noisy input is multiplied to the estimated STFT coefficients by Vocos and then applying inverse STFT to obtain the final waveform which is time-domain normalized and can be directly played. When applying this online normalization, the clip value $\epsilon$ for computing logMel-spectrogram is set to 1e-4 for CleanMel and the input of Vocos generator.
@@ IV Experimental setups
In the section, we present the experiment datasets, experimental configurations, evalution metrics and comparison methods.
@@ IV-A1 Speech enhancement training dataset
The proposed model is trained with synthetic noisy/clean speech pairs. Reverberant speech signals are generated by convolving source speech signals with RIRs, then added with noise signals. Clean speech signals are generated by convolving source speech signals with the direct-path part of RIRs.
In this work we conduct speech enhancement for both Mandarin Chinese and English, and we will evaluate our model on five different datasets (as will be shown later). We attempt to train the model once and test it on all the datasets, which will reflect the general capability of the model under various situations. To do this, we collect data (in terms of source speech, RIRs and noise) from multiple public datasets, and form a training set with sufficient speech quality and environment/device diversity.
Source speech: Source speech signals are collected from 6 datasets, including AISHELL I [28], AISHELL II [29] and THCHS30 [30] for Chinese, and EARS [31], VCTK [32] and DNS I challenge [33] for English. For each language, about 200 hours of high-quality speech data are selected from the original datasets, based on the raw score of DNSMOS P.835 [34]. The selection thresholds are set to 3.6 and 3.5 for Chinese and English data, respectively. Except that the entire EARS training set is included, since EARS involves various emotional speech that cannot be well evaluated by the DNSMOS. The amount of data selected from each dataset is summarized in Table I.
@@
RIR: We use real-measured RIRs from multiple public datasets [35, 36, 37, 38, 39, 40, 41, 42]. Table II shows the statistics of these RIR datasets. For all (even multi-channel) datasets, all RIRs are used, except that we uniformly sampled 1,000 RIRs from the very large original ARNI dataset.
The reverberation time, i.e. $T-{60}$, of RIRs mostly lie in the range of $(0,1.5)$ seconds, except for a few rooms in AIR [36] and NaturalReverb [40]. The distribution of the $T_{60}$s are shown in Fig. 3. Besides the wide distribution range of $T_{60}$s, these RIRs also have large diversity in terms of environments and measuring devices. For example, NaturalReverb [40] is recorded in 271 spaces encountered by humans during daily life, AIR [36] is measured with a dummy head for binaural applications, RWCP [42] use a Head-Torso as source speaker, etc. When synthesizing reverberant speech, 80% source speech samples are convolved with a randomly selected RIR, while there is no RIR convolution for the rest 20% samples to account for the near-field applications where reverberation is negligible.
@@
ACE [35] 84 $[0.33,1.22]$
AIR [36] 204 $[0.08,4.50]$
ARNI [37] 1,000 $[0.51,1.20]$
dEchorate [38] 594 $[0.17,0.75]$
MultiChannel [39] 234 ${0.16,0.36,0.61}$
@@
Noise: Speech and noise are mixed with a random signal-to-noise ratio (SNR) between -5 dB and 20 dB. We use the noise signals from the DNS challenge [43] and the RealMAN dataset [44]. The DNS challenge dataset has about 181 hours of noise sampled from AudioSet [45] and Freesound. The RealMAN dataset has 106 hours of ambient noise recorded in 31 daily life scenes, including various indoor, semi-outdoor, outdoor and transportation scenes.
@@ IV-A2 Speech enhancement evaluation datasets
The speech enhancement performance of the proposed model is evaluated on five datasets. For Chinese, we use the public test set (static speaker) of the RealMAN dataset [44]. For English, the evaluation is conducted on 4 different datasets: (1) the CHiME4 challenge ‘isolated1chtrack’ test set [46], (2) the one-channel test set of REVERB [41], (3) the test set of DNS I challenge [33], (4) the public EARS blind test dataset [31].
IV-A3 ASR evaluation datasets
The ASR performance of the proposed model is evaluated on three datasets. For Chinese, RealMAN (static) [44] is evaluated, with the pre-trained WenetSpeech ASR model provided in ESPNet  [^2] https://github.com/espnet/espnet/tree/master/egs2. For English, CHiME4 [46] and REVERB [41] are evaluated, with their respective pre-trained ASR models obtained using ESPNet. Note that, the ASR models used in this work have the same STFT and mel-frequency configurations, thence one CleanMel model can be used for all of them.
@@ IV-B Configurations
Mode Model size Depth($L+1$) Hidden($H$) #Param(M) FLOPs(G/s)
@@ IV-B1 Data configurations
The sampling rate of all data is set to 16 kHz. STFT is applied using Hanning window with a length of 512 samples (32 ms) and a hop size of 128 and 256 samples (8 and 16 ms) for the offline and online models, respectively. The offline model has a finer temporal resolution than the online model since it is used for ASR in this work and its temporal resolution is aligned with the ASR models. However, we empirically found that, compared to the 16-ms hope size, the 8-ms hope size does not benefit much for the speech enhancement performance. The number of Mel frequencies is set to $F_{\text{Mel}}=80$ for the frequency range of 0-8 kHz. The same STFT implementation (ESPNet implementation [^3] https://github.com/espnet/espnet/blob/master/espnet2/layers/stft.py) is used for the CleanMel networks, neural vocoders and ASR models, to avoid configuration mismatch. The natural logarithm (base of $e$) is used.
@@ IV-B2 Network configurations
Follow [21], the kernel size of the T-Conv1d in the input module and the F-GConv1d layers in the cross-band block are both set to 5. As shown in Fig. 2, in narrow-band block, forward-only and forward/backward Mamba layers are set for online and offline processing, respectively. We set up two model scales for the offline models, referred to as CleanMel-S and CleanMel-L. And the online model scale is set approximately to CleanMel-S. The configurations are shown in Table III. The depth $L$ of online models are set to twice the one of corresponding offline models to have the similar model size. Due to the different setups of STFT hop size, the computational complexity, i.e. FLOPs, of offline models are roughly twice the one of corresponding online models.
@@ IV-B3 Training and inference setups
For training the Vocos neural vocoder [26], we synthesized 400,000 (direct-path) clean speech samples with both English and Chinese data used for CleanMel training. The training configurations keep unchanged as in its original work. For ASR evaluation, as already mentioned, the pre-trained ASR models obtained using in ESPNet are used.
@@ IV-C Evaluation metrics
Speech enhancement performance is evaluated with Perceptual Evaluation of Speech Quality (PESQ) [48], DNSMOS P.808 [49] and P.835 [34], where the background, signal and overall scores for P.835 are all reported. Word Error Rate (WER) and Character Error Rate (CER) are used to evaluate English and Chinese ASR performances, respectively.
@@ IV-D Comparison models
We compare with five advanced speech enhancement models, which were all claimed in their original papers to be able to conduct joint speech denoising and dereverberation, including (i) FullSubNet [15] is a LSTM-based full-band and sub-band fusion network originally proposed for online speech denoising, and extended to speech dereverberation in [16]. For offline processing, we change the uni-directional LSTMs to be bi-directional. (ii) Demucs [50] is an online speech enhancement model that operates directly on waveforms with a U-net. (iii) VoiceFixer [19] also performs Mel-spectrogram enhancement (using a ResUNet) and generates the waveform using a neural vocoder, but is developed only for offline speech enhancement. (iv) StoRM [51] is a powerful diffusion-based offline speech enhancement system. (v) SpatialNet [21] and oSpatialNet [22] perform speech enhancement in the STFT linear-frequency domain, and offer the backbone network for the proposed CleanMel model. In SpatialNet and oSpatialNet, self-attention (and temporal convolution) and Mamba are adopted for learning the narrow-band spatial information in an offline and an online manner, respectively. For offline processing, besides comparing with the original SpatialNet, we also implemented a variant of it with bi-directional Mamba for narrow-band blocks (same architecture with the proposed method), which is referred to as SpatialNet-Mamba, and serves as the linear-frequency baseline for the proposed Mel-frequency model. Note that, SpatialNet and oSpatialNet were originally proposed for multi-channel speech enhancement, and this work is the first one to fully evaluate and analyze the capability of them for single-channel speech enhancement.
@@
To conduct fair comparisons, we re-train the FullSubNet, StoRM and oSpatialNet/SpatialNet models using the same training data utilized for the proposed model. However, we found that it is not easy to re-train the Demucs and Voicefixer models. In [50], different data augmentation strategies are applied when training Demucs on different datasets. Voicefixer [19] was designed for 44.1 kHz signals. Re-training Demucs and VoiceFixer request careful data engineering or hyperparameter search. Therefore, we use the pre-trained checkpoints of Demucs and VoiceFixer to perform speech enhancement on only English data. For Demucs, we use the ‘dns64’ model provided by the authors  [^4] https://github.com/facebookresearch/denoiser. For VoiceFixer, we use the default pre-trained model provided in the open-source package  [^5] https://github.com/haoheliu/voicefixer. Following the VoiceFixer default inference setup, test waveforms are first up-sampled to 44.1 kHz to perform speech enhancement and then down-sampled back to 16 kHz for evaluation.

### v2

@@ III-B2 Cross-band block
The cross-band block learns full-band/cross-band signal dependencies. The original SpatialNet [21] designed this block to learn the linear relationship of inter-channel features (e.g. inter-channel phase difference) across frequencies. In this work, there is no such inter-channel information for the single-channel scenario. Instead, the cross-band block can learn full-band spectral patterns in linear/Mel-frequency domains, which is also critical for (especially single-channel) speech enhancement. The cross-band block processes frames independently along the frequency dimension, with all frames sharing the same network.
Specifically, we retain the cross-band block architecture of SpatialNet, which comprises cascaded frequency convolutional layers (F-GConv1d), across-frequency linear layers (F-Linear), and a second frequency convolutional layer. The frequency convolutional layers perform 1-D convolution along frequency to capture adjacent-frequency correlations, while the across-frequency linear layer processes all frequencies jointly (for each hidden dimension separately) to model full-band dependencies.
One cross-band block is first applied in the linear-frequency domain to learn detailed full-band correlations. To reduce the model complexity, the hidden dimension $H$ is compressed (e.g., to $H/12$) before processing by the across-frequency linear layer ($F^{2}$ complexity for each hidden dimension). After Mel-filterbank transformation, cross-band blocks learn full-band correlations across Mel frequencies, reducing the complexity from $F^{2}$ to $F_{\text{mel}}^{2}$. Correspondingly, we set the hidden dimension to remain $H$ (without compression) to enhance the full-band learning capability.
All the Mel-frequency cross-band blocks share identical across-frequency linear layers.
@@ III-C Back-ends
At inference, CleanMel is followed by either an ASR model or a neural vocoder. Both ASR model and neural vocoder are trained separately from the CleanMel network.
@@ III-C1 ASR
During inference, the enhanced logMel-spectrogram can be directly fed to a pre-trained ASR model, without fine-tuning or joint training. Different ASR systems may use varying STFT configurations, numbers of Mel frequencies, and logarithm bases. To seamlessly integrate the enhanced logMel-spectrogram into one ASR model, CleanMel adopts identical configurations to the target ASR model. Training CleanMel is not very costly, so it can be easily re-trained for a new ASR system, especially for those large-scale or deployed ASR systems. In this work, we only conduct offline ASR combined with offline CleanMel.
@@ III-C2 Neural vocoder
The vocoder we adopt in this work is Vocos [26], a recently proposed Generative Adversarial Network (GAN)-based neural vocoder. The generator of Vocos predicts the STFT coefficients of speech at frame level and then generates waveform through inverse STFT, thence it significantly improves the computational efficiency compared to HiFi-GAN that directly generates waveform at sample level. Vocos uses the multiple-discriminators and multiple-losses proposed in HiFi-GAN [27]. In this work, to unify the front-end and back-end processing, we have made several necessary modifications to Vocos as follows:
  • The magnitude-based Mel-spectrogram of original Vocos is modified as power-based to match front-end and ASR configurations, where the two cases were shown to achieve similar performance in our preliminary experiments.
  • The sampling rate of signals, the STFT configurations and the number of Mel-frequencies of Vocos are adjusted to match the front-end and ASR configurations.
  • The original Vocos is designed for offline processing, as it employs non-causal convolution layers. To enable online processing, we modified Vocos to be causal by substituting each non-causal convolution layer with its causal version. The modified online vocoder still performs well. In addition, to reduce the computational complexity of online processing, the 75% STFT overlap of original Vocos is reduced to be 50% overlap, which still achieves comparable performance.
The offline and online Vocos are used to work with the offline and online CleanMel, respectively. All Vocos models were trained using our direct-path target speech $x(n)$.
@@ III-D Signal Normalization
When cascading separately-trained front-end and back-end models, signal normalization should be performed not only to facilitate the training of respective models but also to align the signal level of cascaded models. Specifically, the level of enhanced logMel of CleanMel should align with the one of the input logMel of ASR and Vocoder models. For offline processing, CleanMel adopts the normalization method of Vocos: a random gain is applied to time-domain signal such that the maximum level of the resulting signal falls between -1 and -6 dBFS (decibels relative to full scale). This ensures the peak sample value approaches but remains less than 1, so that the generated waveform can be directly played at full volume without clipping. CleanMel applies this normalization to the noisy signal, and uses the same gain of noisy signal to the corresponding clean target signal. In this way, the enhanced logMel can be directly fed to Vocos. When applying this time-domain normalization, the logMel clip value $\epsilon$ is set to 1e-5 for both CleanMel and Vocos. ASR models normally have a separate logMel normalization, which will be used to re-normalize the enhanced logMel.
[…]
@@ IV Experimental setups
In the section, we present the experiment datasets, experimental configurations, evaluation metrics and comparison methods.
@@ IV-A1 Speech enhancement training dataset
The proposed model is trained with synthetic noisy/clean speech pairs. Reverberant speech signals are generated by convolving source speech signals with RIRs, then added with noise signals. Clean target speech signals are generated by convolving source speech signals with the direct-path part of RIRs.
In this work, we conduct speech enhancement for both Mandarin Chinese and English. Six different datasets (as will be shown later) are used to evaluate our model. We attempt to train the model once and test it on all the datasets, which will reflect the general capability of the model under various situations. To do this, we collect data (in terms of source speech, RIRs and noise) from multiple public datasets, and form a training set with sufficient speech quality and environment/device diversity.
Source speech: Source speech signals are collected from 6 datasets, including AISHELL I [28], AISHELL II [29] and THCHS30 [30] for Chinese, and EARS [31], VoiceBank [32] and DNS I challenge [33] for English.
For each language, about 200 hours of high-quality speech data are selected from the original datasets, based on the raw score of DNSMOS P.835 [34]. The selection thresholds are set to 3.6 and 3.5 for Chinese and English data, respectively. Except that the entire EARS training set is included, since EARS involves various emotional speech that cannot be well evaluated by the DNSMOS. The two test speakers in the VoiceBank Demand dataset [35] are excluded from our training data.
The amount of data selected from each dataset is summarized in Table I.
@@
RIR: We use real-measured RIRs from multiple public datasets [36, 37, 38, 39, 40, 41, 42, 43]. Table II shows the statistics of these RIR datasets. For all (even multi-channel) datasets, all RIRs are used, except that we uniformly sampled 1,000 RIRs from the very large original ARNI dataset. The reverberation time, i.e. $T_{60}$, of RIRs mostly lie in the range between 0 and 1.5 seconds, except for a few rooms in AIR [37] and NaturalReverb [41]. The distribution of the $T_{60}$s are shown in Fig. 3. Besides the wide distribution range of $T_{60}$s, these RIRs also have large diversity in terms of environments and measuring devices. For example, NaturalReverb [41] is recorded in 271 spaces encountered by humans during daily life, AIR [37] is measured with a dummy head for binaural applications, RWCP [43] use a Head-Torso as source speaker, etc. When synthesizing reverberant speech, 80% source speech samples are convolved with a randomly selected RIR, while there is no RIR convolution for the rest 20% samples to account for the near-field applications where reverberation is negligible.
@@
ACE [36] 84 [0.33,1.22][0.33,1.22][ 0.33 , 1.22 ]
AIR [37] 204 [0.08,4.50][0.08,4.50][ 0.08 , 4.50 ]
ARNI [38] 1,000 [0.51,1.20][0.51,1.20][ 0.51 , 1.20 ]
dEchorate [39] 594 [0.17,0.75][0.17,0.75][ 0.17 , 0.75 ]
MultiChannel [40] 234 0.16,0.36,0.61{0.16,0.36,0.61}0.16 , 0.36 , 0.61
@@
Noise: Speech and noise are mixed at a random signal-to-noise ratio (SNR) between -5 dB and 20 dB. We use the noise signals from the DNS challenge [44] and the RealMAN dataset [45]. The DNS challenge dataset has about 181 hours of noise sampled from AudioSet [46] and Freesound. The RealMAN dataset has 106 hours of ambient noise recorded in 31 daily life scenes, including various indoor, semi-outdoor, outdoor and transportation scenes.
@@ IV-A2 Speech enhancement evaluation datasets
The speech enhancement performance of the proposed model is evaluated on six datasets. For Chinese, we use the public test set (static speaker) of the RealMAN dataset [45]. For English, the evaluation is conducted on 5 different datasets: (1) the CHiME4 challenge ‘isolated_1ch_track’ test set [47], (2) the one-channel test set of REVERB [42], (3) the test set of DNS I challenge [33], (4) the public EARS blind test dataset [31], (5) the test set of VoiceBank Demand dataset [35].
IV-A3 ASR evaluation datasets and models
The ASR performance of the proposed model is evaluated on three datasets. For Chinese, we conduct experiments on the RealMAN (static speaker) dataset [45], using the pre-trained WenetSpeech Conformer ASR model [48] from ESPNet  [^2] https://github.com/espnet/espnet/tree/master/egs2. For English, we evaluate on the CHiME4 [47] and REVERB [42] datasets, using their respective pre-trained E-branchformer [49] and Transformer [50] ASR models from ESPNet, respectively. Note that, the ASR models used in this work have the same STFT and Mel-frequency configurations, thus one CleanMel model can be used for all of them.
@@ IV-B Configurations
[…]
@@ IV-B1 Data configurations
The sampling rate of all data is set to 16 kHz. STFT is applied using Hanning window with a length of 512 samples (32 ms) and a hop size of 128 and 256 samples (8 and 16 ms) for the offline and online models, respectively. The offline model has a finer temporal resolution than the online model since it is used for ASR in this work and its temporal resolution is aligned with the ASR models. However, we empirically found that, compared to the 16-ms hop size, the 8-ms hop size does not benefit much for the speech enhancement performance. The number of Mel frequencies is set to $F_{\text{mel}}=80$ for the frequency range of 0-8 kHz. The same STFT implementation (ESPNet implementation [^3] https://github.com/espnet/espnet/blob/master/espnet2/layers/stft.py) is used for the CleanMel networks, neural vocoders and ASR models, to avoid configuration mismatch. The natural logarithm (base of $e$) is used.
@@ IV-B2 Network configurations
Follow [21], the kernel size of the T-Conv1d in the input module and the F-GConv1d layers in the cross-band block are both set to 5. As shown in Fig. 2, in narrow-band block, forward-only and forward/backward Mamba layers are set for online and offline processing, respectively. We set up two model scales for the offline models, referred to as CleanMel-S and CleanMel-L. The online model scale is set approximately to CleanMel-S. The configurations are shown in Table III. The depth $L$ of online models are set to twice the one of corresponding offline models to have the similar model size. Due to the different setups of STFT hop size, the computational complexity, i.e. FLOPs, of offline models are roughly twice the one of corresponding online models.
@@ IV-B3 Training and inference setups
For training the Vocos neural vocoder [26], we synthesized 400,000 (direct-path) clean speech samples with both English and Chinese data used for CleanMel training. The training configurations remain unchanged as in its original work.
@@ IV-C Evaluation metrics
Speech enhancement performance is evaluated with Perceptual Evaluation of Speech Quality (PESQ) [52], DNSMOS P.808 [53] and P.835 [34], where the background (BAK), signal (SIG) and overall (OVL) scores for P.835 are all reported. Word Error Rate (WER) and Character Error Rate (CER) are used to evaluate English and Chinese ASR performances, respectively.
@@ IV-D Comparison models
We compare with advanced speech enhancement models, most of which were claimed in their original papers to be able to conduct joint speech denoising and dereverberation, including (i) ConvTasNet [1] is a fully convolutional time-domain (offline) speech separation network. (ii) FullSubNet [15] is a LSTM-based full-band and sub-band fusion network originally proposed for online speech denoising, and extended to speech dereverberation in [16]. For offline processing, we change the uni-directional LSTMs to be bi-directional. (iii) Demucs [2] is an online time-domain speech enhancement network with a U-net architecture. (iv) CMGAN [54] is a two-stage Conformer-based offline speech enhancement network. (v) LiSenNet [55] is a super light-weight online speech enhancement network primirily with a dual-path recurrent module. Both CMGAN and LiSenNet utilize mixed training losses, including speech regression losses and an adversarial loss. (vi) VoiceFixer [19] also enhances Mel-spectrograms (using a ResUNet) and generates the waveform using a neural vocoder, but supports only offline processing. (vii) CDiffuse [56], StoRM [57] and UNIVERSE++ [58] are three diffusion-based offline speech enhancement models. (viii) SpatialNet [21] and oSpatialNet [22] perform speech enhancement in the STFT linear-frequency domain, and offer the backbone network for the proposed CleanMel model. In SpatialNet and oSpatialNet, self-attention (and temporal convolution) and Mamba are adopted for learning the narrow-band spatial information in an offline and an online manner, respectively.
For offline processing, besides comparing with the original SpatialNet, we also implemented a variant of it with bi-directional Mamba for narrow-band blocks (same architecture with the proposed method), which is referred to as SpatialNet-Mamba, and serves as the linear-frequency baseline for the proposed Mel-frequency model. Note that, SpatialNet and oSpatialNet were originally proposed for multi-channel speech enhancement, and this work is the first one to fully evaluate and analyze the capability of them for single-channel speech enhancement.
@@
To conduct fair comparisons, we re-trained the ConvTasNet, FullSubNet, CMGAN, LiSenNet, CDiffuse, StoRM, UNIVERSE++ and oSpatialNet/SpatialNet models, using the same training data utilized for the proposed model. LiSenNet was trained without the noise detector [55]. UNIVERSE++ was trained under its universal speech enhancement configuration [58]. All models were trained to convergence according to their model selection strategy. However, we found that it is not easy to re-train the Demucs and Voicefixer models. In [2], different data augmentation strategies are applied when training Demucs on different datasets. Voicefixer [19] was designed for 44.1 kHz signals. Re-training Demucs and VoiceFixer requires careful data engineering or hyperparameter search. Therefore, we used the pre-trained checkpoints of Demucs and VoiceFixer to perform speech enhancement on only English data. For Demucs, we used the ‘dns64’ model provided by the authors  [^4] https://github.com/facebookresearch/denoiser. For VoiceFixer, we empolyed the default pre-trained model provided in the open-source package  [^5] https://github.com/haoheliu/voicefixer. Following the VoiceFixer default inference setup, test waveforms were first up-sampled to 44.1 kHz to perform speech enhancement and then down-sampled back to 16 kHz for evaluation.